OpenStudy (anonymous):

A model airplane is shot up from a platform 1 foot above the ground with an initial upward velocity of 56 feet per second. The hight of the airplane above ground after t seconds is given by the equation h=-16t^2+56t+1 where h is the hight of the airplane in feet and t is the time in seconds after it is launched. Approximately how long does it take the airplane to reach its maximum height? A)0.3 seconds B)1.8 seconds C)3.5 seconds D)6.9 seconds Please Help

OpenStudy (anonymous):

what jerk would shoot down a model airplane! :( It's my favorite toy!

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

haha an idiot who hates aviation @sourwing I'm a pilot in training so I personally dislike it lol

OpenStudy (anonymous):

dont you have a graphing calculator?

OpenStudy (anonymous):

yeah but I just don't get it period can you give an explanation @sourwing ?

OpenStudy (anonymous):

oh wait, it asks for maximum height -.- well, the vertex any parabola is at x = -b/(2a)

OpenStudy (anonymous):

vertex *of* any parabola

OpenStudy (anonymous):

t = -56 / (2 * -16) <-- ans

OpenStudy (anonymous):

this is trigonometry, i believe i am learning about this now. that is a triangle are formula. i would draw it and find where x is located. Then you have to round off the answer. it would be 1.8 (-56 \0/(2*-16) = 1.75 = 1.8

OpenStudy (anonymous):

thanks both of you!!